OpenStudy (anonymous):

A golf ball is hit from the ground with an initial velocity of 208 feet per second. Assume the starting height of the ball is 0 feet. How long will it take the golf ball to hit the ground? 13 secs 21 secs 42 secs 6 secs

OpenStudy (anonymous):

13 lol i am in the middle of doing my work too :D

OpenStudy (anonymous):

flvs?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

can you help with this one

OpenStudy (anonymous):

Use complete sentences to explain how you would use graphing technology to find the solutions of 0 = −x2 + 2.5x + 1. What do the solutions of 0 = −x2 + 2.5x + 1 represent on the graph of the equation?

OpenStudy (anonymous):

im stuck on that 1 too! FML

OpenStudy (anonymous):

oh thats fine

OpenStudy (anonymous):

if i find any i will make sure to say

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@JasonM3596 i solved it myself on geogebra but thanks for the help

OpenStudy (anonymous):

@JasonM3596 if you use geogebra it is easy and explains itself

OpenStudy (anonymous):

whats the answer :P

OpenStudy (anonymous):

basically: you just start geogebra then type the equation into geogebra and you see a line that looks like it bends around in a circle. you will have to rephrase a little or it will be plagiarizing

OpenStudy (anonymous):

i will get it! XD